The octopus above is swimming along when it decelerates at 1.5ms-2 for 0.2s to a final speed of 3ms-1. What is the octopus's initial speed?
1
Expert's answer
2019-08-08T10:54:12-0400
"a = {v - v_0 \over t}"

"v_0 = v - at"

Octopus decelerates, therefore a = -1.5 m/s2. In the end:

"v_0 = 3 - (-1,5)*0.2 = 3 + 0.3 = 3.3 m\/s"

Answer: 3.3 m/s.